What is the volume of CO2 at STP formed by the combustion of 290 g of butane ?

Grade:10

1 Answers

Arun
5 years ago

C4H10 + 6.5 O2= 4CO2 + 5 H2O

So, four moles of CO2 for each mole of Butane.

Molar mass of Butane = (4x12) + (10 x 1) = 58g/mol

Moles of Butane = 290/58 = 5 mol

So, if we're assuming 24 L per mole of gas then 120 L of Butane.

Which gives 480 L ofCO2

Although other molar gas volumes are available as it is temperature dependent